### v2

# A systematic review of challenges and proposed solutions in modeling multimodal data

Abstract
Multimodal data modeling has emerged as a powerful approach in clinical research, enabling the integration of diverse data types such as imaging, genomics, wearable sensors, and electronic health records. Despite its potential to improve diagnostic accuracy and support personalized care, modeling such heterogeneous data presents significant technical challenges. This systematic review synthesizes findings from 69 studies to identify common obstacles, including missing modalities, limited sample sizes, dimensionality imbalance, interpretability issues, and finding the optimal fusion techniques. We highlight recent methodological advances, such as transfer learning, generative models, attention mechanisms, and neural architecture search that offer promising solutions. By mapping current trends and innovations, this review provides a comprehensive overview of the field and offers practical insights to guide future research and development in multimodal modeling for medical applications.
keywords:
Multimodal data, challenges, Fusion strategy, Modeling
1 Introduction
In healthcare, data comes in various modalities like electronic health records (EHR), imaging, genomics, and wearable devices. Integrating more than one of these modalities in a multimodal data analysis holds significant promise for improving our understanding and management of complex health conditions 1. One notable domain where multimodal data analysis can make a transformative impact is the management of chronic diseases like asthma and chronic obstructive pulmonary disease (COPD) patients 2. Here, medical imaging data, such as computerized tomography, provides insights into lung morphology and pathological changes. Wearable devices monitor physiological parameters and activity levels, offering real-time data on the patient’s disease trajectory. Environmental data, encompassing air quality and pollutants, helps identify external factors influencing respiratory conditions. Clinical records and genomic information contribute to personalized treatment strategies, while patient-reported outcomes capture subjective experiences, shaping a patient-centric approach. Combining several or even all of these modalities allows us to understand the disease and reveal interactions between different aspects.
Modeling multimodal data is a challenging task that extends well beyond the analysis of individual modalities. Specifically, each modality may have different data types, formats, scales, levels of noise, and levels of interpretability. Moreover, modalities may exhibit varying degrees of completeness, sparsity, and dimensionality, making it challenging to integrate and analyze them together. In context of modeling multimodal data, fusion techniques are of great importance. A fusion strategy involves combining and integrating the different modalities to create a unified representation that captures the complementary information from each modality. These techniques can be divided into four main categories: (i) early fusion, (ii) intermediate fusion, (iii) late fusion, and (iv) hybrid fusion. (i) Early fusion involves integrating information from different sources at the data level without learning any marginal representations or feature extraction, treating them as unimodal input. Although early fusion offers simplicity and the ability to learn cross-modal relationships, it may pose challenges when dealing with different dimensionalities and sampling rates across modalities. Moreover, it overlooks the relationships between different modalities and may not be suitable for heterogeneous modalities, such as combining image data with omics data. (ii) Intermediate fusion, in contrast, involves learning modality-specific representations to identify correlations within each modality before generating joint representations or making predictions. This approach offers flexibility in selecting different architectures for feature extraction from various modalities 3. By employing distinct architectures and layers for each modality, it can effectively handle heterogeneous data. Additionally, intermediate fusion is resilient to dimensionality imbalances 3 and can accommodate missing modalities without significantly affecting performance 3, 4. (iii) Late fusion, on the other hand, trains separate models for each modality and integrates their individual predictions at a later stage. This method is particularly well-suited for managing heterogeneous modalities and addressing dimensionality imbalances and missing data. However, it may be less effective for modalities with strong correlations, as it does not capture interactions between different data sources. In addition, in scenarios where specific information holds greater significance, using late fusion offers notable benefits. The model prioritizes the most important information it has, then adds more details from other sources to improve its understanding. This prioritization allows for a more nuanced integration of information, enhancing the model’s ability to make informed decisions and predictions. Additionally, in scenarios where certain information is more critical, late fusion provides significant advantages. The model first prioritizes the most relevant information, then incorporates additional details from other sources to refine its understanding. This approach enables a more nuanced integration of data, improving the model’s ability to make informed decisions and predictions. (iv) Hybrid fusion combines elements from the aforementioned methods, allowing for flexible processing across multiple stages. This approach is especially effective in scenarios with diverse data modalities, where early fusion can be used for closely related data types, and later fusion stages facilitate the integration of more distinct modalities. While less frequently applied than the three primary fusion categories, hybrid fusion provides the adaptability needed to meet the specific requirements of multimodal models. However, selecting the optimal fusion technique represents only one of the many challenges faced in multimodal modeling 5. The integration of diverse data modalities into a multimodal model is visually represented in Figure 1. The figure visually represents the three primary fusion strategies (early, intermediate, and late), for simplicity, it excludes hybrid and other less frequently used strategies.
Figure 1: Multimodal Modeling: Various information sources can be collected from a patient during a clinical stay or a doctor’s visit resulting in different data modalities. Subsequently, the modalities can be evaluated jointly using a suitable fusion technique, where the fusion (F) can be happening at different stages in the process.
The objective of this systematic review is to identify the challenges encountered in modeling multimodal data in medical domains. We aim to provide insight into the key challenges and their underlying causes, as well as potential solutions to address these challenges.
Several review articles have discussed the application of deep learning to analyze multimodal data. For instance, Mohsen et al. 6 explored integrating multiple data modalities with artificial intelligence models for a variety of applications, including combining medical imaging with data from electronic health records (EHRs). Similarly, Azam et al. 7 and Zhang et al. 8 solely focused on fusing different imaging modalities. Behrad and Abadeh 9 and Stahlschmidt et al. 3 primarily concentrated on integrating omics data with other modalities. Huang et al. 10 address the fusion of multimodal EHR data using traditional machine learning and deep learning methods. Pawłowski et al. 11 and Shaik et al. 12 focus on fusion techniques and feature extraction methods.
While these reviews provide important insights into various fusion strategies, our work differs in its broader examination of technical challenges across diverse medical domains, along with a comprehensive exploration of solutions tailored to the clinical context. We aim to fill this gap by gathering the latest advancements in deep learning approaches for representing multimodal data and summarizing various research efforts in terms of the challenges they confront and the techniques they employ to overcome them. Although this systematic review focuses specifically on the challenges associated with multimodal data in the medical domains, some references included in the manuscript are not part of the systematic review. These additional references have been included to provide essential background information. By incorporating these sources, we aim to offer a more comprehensive context that complements the findings of the review.
2 Results
The initial search yielded a total of 801 publications obtained from PubMed after removing duplicates. Following the title and abstract screening of these 801 papers, we identified 84 papers that met our initial inclusion criteria. After thorough evaluation, we retained 58 papers that specifically addressed challenges related to the modeling of multimodal data in the medical domains as outlined in Figure 2. Papers that did not directly focus on these modeling challenges or dealt with other domains such as clinical challenges were excluded. In addition, we manually searched Google Scholar and included 11 more relevant papers, increasing the total to 69 papers.
Figure 2: PRISMA flow diagram illustrating the study selection process following the systematic search.
We compiled our findings into a summary table to provide a comprehensive overview of the fusion strategies within the literature (see Table 1).
Table 1: Summary of the 69 included studies categorized by fusion strategies for highlighting the distribution of fusion approaches.
Fusion Strategy Number of Studies
Intermediate fusion 24
Hybrid fusion 15
Late fusion 14
Early fusion 7
Hierarchical fusion 2
Probability-based fusion 1
Maximum correlation-based fusion 1
Not indicated 5
Intermediate fusion emerged as the most widely used approach, applied in 24 studies. Hybrid fusion was employed in 15 studies, while late fusion and early fusion strategies were utilized in 14 and 7 studies, respectively. Two studies employed hierarchical fusion strategies to tackle challenges in multimodal data integration 13, 14. Zhang et al. 14 introduced a transformer-based model that utilizes hierarchical multimodal transformers to progressively integrate features from shallow to deep layers. This design facilitates effective cross-modality interaction and addresses issues such as spatial misalignment and data heterogeneity, distinguishing it from traditional early, intermediate, or late fusion techniques. Similarly, Li et al. 13 proposed a hierarchical factorized bilinear fusion strategy tailored for cancer survival prediction, which integrates genomic and pathological image data across low- and high-level fusion stages. While both approaches highlight the critical role of cross-modality interactions, this method further incorporates modality-specific and cross-modality attentional factorized bilinear modules, enabling reduced computational complexity and enhanced performance, particularly in genomic data analysis. Together, these studies underscore the adaptability and effectiveness of hierarchical fusion frameworks in optimizing multimodal data integration across diverse applications. One study 15 employed a probability-based fusion approach, which involves calculating the probability of each feature being relevant and discriminative for classification, as determined by two optimization algorithms. Features with higher probabilities are prioritized and included in the final fused feature set, ensuring the combination of only the most informative and non-redundant features for classification. One other study 16 used the maximum correlation-based fusion approach, which emphasizes integrating features with the strongest cross-modal correlations. In this method, features from different modalities are analyzed for their pairwise correlations. Strongly correlated features are selected and combined, while weakly correlated or redundant features are excluded. This strategy enhances the representational power of the fused data by focusing on meaningful relationships while reducing noise and redundancy. Finally, five studies did not specify or clearly define their fusion strategy.
Table 2 provides an overview of the number of publications across different medical domains, along with the distribution of modeling challenges within these domains. This classification highlights the primary areas of application for multimodal data modeling, with neurology emerging as the most frequently studied domain (22 studies), followed by oncology (20 studies), physiology (6 studies), and ophthalmology and psychiatry (3 studies each). Additional domains such as psychology, endocrinology, cardiology, orthopedics, pharmacology, and gastroenterology further demonstrate the adaptability of multimodal data modeling across various medical contexts. Additionally, three studies explored non-medical applications focusing on synthetic data. Notably, neurology and oncology also stand out as the domains facing the highest number of modeling challenges, particularly regarding missing modalities, small datasets, and the identification of optimal fusion techniques.
The yearly frequencies of the identified challenges are illustrated in Figure Figure 3. Specifically since 2020, there was a strong increase in studies addressing challenges in modeling multimodal data (the number of studies in 2023 is lower as the search strategy was limited until October 2023). This upward trend indicates that research in this area will likely continue to expand, addressing new challenges as multimodal applications become more widespread.
Figure 3: Histogram of key challenges addressed in yearly publications. Studies may address multiple challenges.
Table 2: Overview of modeling challenges reported across different medical domains in the included studies. Each row corresponds to a medical domain and summarizes the number of studies (second column) and the frequency of each identified challenge in that domain. Note that a single study may address multiple challenges; hence, the values in each row may sum to more than the total number of studies. A dash (-) indicates that no studies reported that specific challenge in the given domain.
Medical Domain Number of Missing Interpretability Small Imbalance Optimal Other
Studies Data Data Dimensionality Fusion
Neurology 22 8 2 4 2 4 8
Oncology 20 5 3 6 2 4 8
Physiology 6 2 - - 2 1 2
Ophthalmology 3 0 1 3 1 - -
Psychiatry 3 - 1 - - 1 2
Dermatology 2 - 1 1 - - -
Endocrinology 2 - 1 2 - 1 -
Intensive care 2 - 1 - - - 2
Pneumology 2 - 2 - - - -
Cardiology 1 - - - - - 1
Gastroenterology 1 - - - - 1 -
Orthopedics 1 - 1 - - - -
Pharmacology 1 - 1 - - - 1
Non medical 3 - - 1 - 1 1
LABEL:Tab:challenges highlights for each study the particular challenges they tackled, along with the strategies they implemented to address these effectively. They are discussed in more detail and according to the challenges in the following section. Further explanation of the proposed solutions for each challenge, including the medical domain, disease type and fusion strategy used in each publication, can be found in the Supplementary file.
Table 3: Summary of challenges and solutions identified in 69 studies: The challenges are indicated by colored boxes   Missing modality,   Small data,   Interpretability,   Imbalance in dimensionality,   Optimal fusion,   Other, and the solutions to address the challenges are written next to the boxes. Some of the publications addressed more than one challenge, and some publications addressed multiple challenges with the same method. Additionally, for “other” challenges that fell into different categories, we first stated the specific challenge followed by a colon and then presented the corresponding solution.
Study publication Challenges & Solutions
Cai et al. 17   Frequency Domain Transformer model, which leverages multimodal data (histopathological images and genomic data)
Cremonesi et al. 18   Interoperability of Data for federated learning: Common Data Model and standardizing data representation and semantics across the different sources of data
Dai et al. 19   Vision Transformer to fuse multi-modal features by capturing interactions and long-range dependencies   Loss of interaction between modalities: joint attention module based on the vision Transformer
Hou et al. 20   Incorporating an online masked autoencoder into the hybrid graph convolutional network
Krix et al. 21   Using the integrated gradients method to obtain feature importance and attention mechanisms to make predictions explainable   Capturing complex relationships: employing graph convolutional networks and graph attention networks
Li et al. 22   Gradient-weighted class activation mapping method   Transfer learning   Focal loss function
Martí-Juan et al. 23   Variational autoencoder
Ning et al. 24   Latent representation learning
Partin et al. 25   Data augmentation   Weighted loss function
Rahaman et al. 26   Mid-fusion approach with an attention mechanism   Saliency analysis is applied to the trained model to identify the most important features from each modality
Shao et al. 27   Discovering significant features: Metric learning, Hilbert-Schmidt Independence Criteria
Shi et al. 28   Feature selection: Generative adversarial network
Sükei et al. 29   Hidden markov model and including information from unlabelled samples via transfer learning
Wang et al. 30   Modeling the fully connected layer of a deep neural network as a potential function following the classical Gibbs measure
Xia et al. 31   Introducing an adaptive fusion layer that assigns weights to different modalities based on their importance
Xing et al. 32   Gradient-guided knowledge modulation scheme
Zhang et al. 14   Throughout fusion transformer
Aghili et al. 33   Gradient boosting
Ahuja et al. 34   Guided topic priors by leveraging prior information from informative surrogate features   Scalability: time-efficient variational Bayesian inference implemented with parallelization in C++
Boehm et al. 35   Synthesizing complementary modalities of data with emerging multimodal artificial intelligence methods
Cai et al. 36   Vision Transformer
Golovanevsky et al. 37   Capturing interactions between modalities: Cross-modal bi-directional attention
Hatami et al. 38   Weighted Ensenble Fusion
Jabeen et al. 15   Data augmentation, transfer learning   Features selection using the differential evolution and introducing probability-based fusion method
Karaman et al. 39   Re-weighted loss function
Li et al. 40   Identifying key features contributing to predictions through feature importance ranking methods
Li et al. 13   Limited depth in multimodal information mining: Hirarchical fusion
Lim et al. 41   Transfer learning
Rabinovici-Cohen et al. 42   SHapley Additive exPlanations   Hetrogeneity: Stratification
Saad et al. 43   Deep convolutional neural network   Multi-task learning, Synthetic data generation   Heterogeneity: Multimodal feature fusion
Sheu et al. 44   Hybrid model using SHAP and deep learning based on dense layers-batch normalization
Vanguri et al. 45   Dynamic deep attention-based multiple-instance learning model that incorporates a masking mechanism   This masking mechanism dynamically adjusts attention weights to ignore missing modalities while effectively using the available data to support robust prediction
Wang et al. 46   Gradient-weighted class activation mapping, Shapley Additive exPlanations
Wang et al. 47   knowledge-infused global-local data fusion model
Xu et al. 48   Shapley additive explanations   Variational autoencoder
Xu et al. 49   Deep latent representation sequence learning   Recurrent neural network
Yoo et al. 50   Transfer learning, Gradient Boosting
Ziegler et al. 51   Multimodal conditional generative adversarial networks
Abdelaziz et al. 52   Linear interpolation   Heterogeneity: Convolutional neural network
An et al. 53   Capturing the complex relationships and temporal dynamics among different modalities and account for the irregularities in patient visit timings: Proposing Time-Aware Multi-Type Data Fusion Representation Learning: Uses bidirectional long short-term memory and convolutional neural network networks to extract temporal and spatial patterns from different data types and applies a self-attention mechanism to combine and highlight key features across data types.
Bahador et al. 54   High-dimensional data and intensive computational requirements: low-level fusion technique
Dong et al. 55   Hypergraph laplacian regularization term that captures high-order relationships among subjects   Heterogeneity: high-order Laplacian regularized low-rank representation
Guan et al. 56   Knowledge distillation
Kustowski et al. 57   Transfer learning
Mao et al. 58   Focal loss function   Capturing intermodality interaction: Convolutional neural network, cross-modality interaction module
Shickel et al. 59   Transfer learning
Xin et al. 60   Interpretable Deep Multimodal Fusion Framework based on Deep Canonical Correlation Analysis
Bi et al. 61   Proposing the Clustering Evolutionary Random Forest to process high-dimensional data by clustering decision trees in a random forest and retaining only the most accurate ones.
Brand et al. 62   Temporal dependencies: integrating temporal relationships using $\ell_{2,1}$-norm regularization to ensure that features are consistent over time, enabling the identification of longitudinal biomarkers
Khan et al. 16   Features selection using the differential evolution and introducing maximum correlation-based fusion approach
Li et al. 63   Heterogeneity: MixEHR, two-level probabilistic joint matrix factorization approach
Liu et al. 64   Using multi-task deep learning models   Combining deep learning scores with radiologist-assessed imaging features and clinical factors in Cox models for transparent survival prediction   Heterogeneity: Multi-task learning
Liu et al. 65   High Dimensionality: Multi-Task Feature Selection to jointly learn feature selection across modalities and remove irrelevant/redundant features
von Lühmann et al. 66   General linear model with temporally embedded Canonical Correlation Analysis
Wang et al. 67   Knowledge distillation
Hao et al. 68   Proposing a biologically interpretable pathway-based sparse deep neural network called Cox-PASN
Tiulpin et al. 69   Using Gradient-weighted Class Activation Mapping
Alam et al. 70   Extracting high order interactions within and among data modalities: Introducing a semiparametric kernel machine approach, that captures pairwise and higher-order Hadamard product interactions across features from different modalities
Cai et al. 71   Multimodal generative adversarial networks
Vivar et al. 72   Matrix completion
Zille et al. 73   Extracting relationship between different modalities: Combining Lasso regression and sparse canonical correlation analysis in a collaborative regression framework to enforce cross-correlation in regression problems
Lei et al. 74   Dimension reduction/Feature selection
Tran and Liatsis 75   Combining one-class classification and Real Adaboost
Eshaghi et al. 76   Kernel-Based   Heterogeneity: Kernel-Based
Liu et al. 77   Multiple kernel learning, Group Lasso regularization   Heterogeneity: Multiple kernel learning
Zhang et al. 78   High-order correlation among features: Feature Correlation Hypergraph
Thung et al. 79   Low-rank matrix completion algorithm
Madabhushi et al. 80   Based on non-linear dimensionality reduction methods   Heterogeneity: Kernel-Based
Tiwari et al. 81   Semi-Supervised Multi-Kernel Graph Embedding
2.1 Challenges and solutions
2.1.1 Missing of entire modality
In multimodal data settings, there are typically two types of missing data: the first involves missing values within a modality, where individual modalities contain incomplete information, and the second involves the complete absence of an entire modality for certain samples. This means that for some data points, one type of information might be entirely missing, adding an extra layer of complexity. Addressing this challenge requires specialized imputation approaches distinct from those used for handling missing values within individual modalities. Missing modalities are a frequent issue in medical settings, as various examinations are typically conducted for different patients. Moreover, patients may lack access to certain modalities due to factors such as high costs, safety concerns, sensor malfunction, data corruption, or patient dropout 82.
While this systematic review focuses on missing modalities in multimodal data, it’s important to acknowledge the broader literature on missing data, which offers foundational approaches for handling such challenges. Traditional methods for handling missing data involve simple imputation methods 83, which replace missing values with mean or median from the known data, and model-based imputation 84 which uses multivariate statistical methods and leverages feature dependencies to impute missing values. Examples of such model-based methods include k-nearest neighbors 85, MissForest (which uses Random Forests) 86, and matrix completion methods 87. Additionally, multiple imputation methods 88 create several different plausible datasets by filling in missing values multiple times and combining the results to account for the uncertainty of the imputed values. However, in the context of multimodal datasets, these approaches face limitations. Multimodal datasets contain different types of information from various sources for the same subject, and these approaches do not effectively consider the correlations between these modalities (essential information missed in one modality can be maintained in another modality). For instance, simple imputation methods might oversimplify complex relationships present between modalities, model-based imputation relies on assumptions of linear relationships among modalities, and deep learning methods might not explicitly leverage cross-modality correlations. Since each modality contributes unique insights that need to be considered collectively for accurate imputation, it is necessary to develop techniques that can effectively capture and leverage these interdependencies between different modalities.
Common deep learning-based approaches, such as Variational Autoencoder (VAE)-based and Generative Adversarial Network (GAN)-based methods, provide innovative solutions for missing data imputation. However, they face challenges in capturing multimodal correlations. For instance, the VAE model proposed by McCoy et al. 89 can handle missing data but does not account for cross-modality relationships. To address this, Wu and Goodman 90 extended the VAE framework to the Multimodal Variational Autoencoder (MVAE), which utilizes a product of expert inference networks to better integrate unique modality contributions 91. However, MVAE struggles to handle randomly missing modalities effectively due to its reliance on fixed input-output structures, limiting its use in clinical applications.
Several studies from our systematic review highlight the limitations of commonly used imputation techniques in multimodal data settings and propose more advanced methods to address this challenge. In total, 15 publications in this review tackled the issue of missing modalities, underscoring its significance in multimodal modeling.
For taking the correlation between modalities into account, Saad et al. 43 proposed using a multitasking learning approach. This approach involves dividing incomplete multimodal datasets into subsets, each containing a different combination of modalities. Individual classifiers are then trained independently on these subsets, and their outputs are combined to generate the final classifier. However, this method assumes linear relationships between modalities, which is often unrealistic. Another approach, proposed by Vivar et al. 72, involves matrix completion techniques by concatenating multimodal features into a single matrix. Yet, this approach fails when entire modalities are missing, as missing entries form blocks rather than random distributions, violating assumptions of matrix completion methods. To address these limitations, extensions of VAE and GAN frameworks have been proposed. For example, Xu et al. 48 introduced the Dynamic Multimodal Variational Autoencoder (DMVAE), which incorporates modality indicators for missing modalities and leverages joint representations to capture correlations across modalities. This enables DMVAE to impute missing features more effectively than MVAE. However, DMVAE is not well-suited for image data. In this context, GAN-based approaches provide alternative solutions. Cai et al. 71 adapted GANs for multimodal settings, using an encoder-decoder architecture to generate missing modalities based on existing ones. To extend this capability beyond imaging, Ziegler et al. 51 proposed the Multimodal Conditional GAN (MMCGAN). This model combines conditional GANs for tabular data with models for conditional 3D image synthesis, preserving cross-modal correlations and generating realistic synthetic data across diverse modalities. In addition to these generative methods, Hou et al. 20 introduced a Hybrid Graph Convolutional Network that incorporates an online masked autoencoder to leverage cross-modal dependencies via transformers. However Wang et al. 67 highlighted a drawback of imputation-based methods, noting that they can introduce noise, especially when only a limited number of complete modality samples are available, which can negatively affect downstream tasks. To address this, they proposed a novel multimodal learning framework that leverages all available samples, including those with incomplete modalities, without relying on imputation. Their approach is based on knowledge distillation 92. In this framework, modality-specific models are first trained independently on all available data, even if some modalities are missing. These models act as “teachers” to train a multimodal “student” model. The student model learns by combining the soft labels provided by the teacher models with the true one-hot labels. Because each teacher model is trained exclusively on a single modality, it benefits from a larger effective sample size, making it a robust expert for its respective modality. The student model integrates these modality-specific insights, enabling it to combine information across all modalities effectively. This approach stands out because it does not discard incomplete modality samples or directly impute missing data. Instead, it ensures that even incomplete samples contribute to the expertise of the teacher models, maximizing data utilization 67. However, Du et al. 93 noted that not all knowledge sources are equally reliable, and conflicting information can arise during distillation. Simply weighing all knowledge sources equally can reduce the effectiveness of the student model. To address this limitation, Xing et al. 32 proposed the Gradient-guided Knowledge Modulation (GKM) scheme. GKM dynamically adjusts the influence of each knowledge source based on its consistency in the gradient space. By prioritizing reliable information and filtering out misleading knowledge, GKM enhances the distillation process, ensuring the student model absorbs useful insights while minimizing conflicts. This selective adjustment improves the robustness and performance of the student network, offering more dependable guidance from diverse sources.
2.1.2 Small dataset
One of the major challenges in multimodal data fusion is dealing with small data. In total, 17 publications in this review tackled this challenge, reflecting its critical importance in multimodal modeling Limited, insufficient, or sparse data in multimodal fusion complicates the integration of multiple modalities, making it challenging to capture the true underlying patterns and relationships. Researchers have proposed various innovative approaches to address the small data challenge, ranging from data augmentation to transfer learning. Data augmentation techniques 15 stand out, artificially expanding dataset size and diversity, thereby enhancing the training of robust fusion models. Another strategy involves multimodal pre-training, where models learn generic representations from large unimodal datasets, applicable to smaller multimodal datasets. Multi-task learning 94, 64 emerges as a powerful technique, enabling the model to jointly train on related tasks, leveraging shared information across modalities to improve performance, even with limited data. Ensemble learning, combining multiple models or techniques, also proves effective in addressing the small data challenge. On the other hand, Cai et al. 36 proposed a novel multimodal transformer for image and metadata fusion, incorporating pre-trained Vision Transformer (ViT) models.
In the context of medical imaging, Li et al. 22 proposes transfer learning methods, involving fine-tuning CNN models pre-trained on natural image datasets to overcome the scarcity of medical image data. Similarly, Shickel et al. 59, Lim et al. 41, Jabeen et al. 15 and Yoo et al. 50 overcome small sample size challenge through transfer learning. This involves initializing a smaller model by leveraging pre-existing models trained on larger datasets. These initialized models are subsequently fine-tuned using the smaller, domain-specific dataset, thereby optimizing performance and generalization for medical image analysis tasks
However, Lei et al. 74 argued that the “small-n-large-p” problem, characterized by a large feature dimension and small sample size, presents persistent challenges in identifying clinical subjects correctly via robust modeling. Both dimension reduction and feature selection offer promising solutions, addressing over-fitting issues and reducing computational time. Saad et al. 43, taking a different approach to tackle this challenge, they utilize traditional canonical correlation analysis for data fusion highlighting its efficacy in leveraging hidden linear correlation among features that are both small size and heterogeneous. Which can overcome the challenge of small datasets, particularly for 2D images and 1D genomic data.
Kustowski et al. 57 argued against a purely data-driven approach with limited data and proposed a hybrid methodology combining knowledge-driven and data-driven approaches. The process begins with a simulation-trained neural network surrogate model, which captures crucial correlations among different data modalities and between simulation inputs and outputs. Subsequently, this knowledge is transferred to retrain the model, aligning it with real-world data for improved predictive performance. In another approach, Guan et al. 56 addressed the challenge of limited data by using knowledge distillation. This technique transfers insights from a more complex teacher network to a smaller student network 95. In a similar way, the knowledge obtained from multimodal data by a teacher model can be transferred to a student model. Faced with a scarcity of Magnetic Resonance Imaging (MRI) data, Guan et al. 56 propose a multi-modal multi-instance distillation strategy. This approach aims to transfer insights from a multi-modal teacher network to a single-modal student network. The teacher network, trained with both MRI and clinical data, captured comprehensive disease-related patterns. This knowledge was distilled into the student network, which was trained only on MRI data, ensuring that the richer, multi-modal insights were preserved and utilized effectively.
2.1.3 Interpretability
Interpretability in multimodal modeling refers to understanding how a model integrates and assigns significance to information from various sources, making its predictions explainable. This is particularly important in medical fields, where decisions made based on model outputs can directly impact patient health. In this review, 14 publications address the challenge of interpretability, highlighting its essential role in supporting model reliability and understanding in healthcare applications. Various techniques have been developed to improve the interpretability of deep learning models in multimodal environments. One such method is the use of attention mechanisms, which allow the model to focus on specific aspects of the data during prediction. For instance, these mechanisms can highlight important regions in a medical image or key elements in genomic sequences, thereby improving transparency in the prediction process. Another approach involves gradient-based attribution algorithms, which identify and quantify the contribution of each modality to the final prediction. These methods are particularly useful in understanding the biological relevance of the model’s outputs. For instance, to increase the interpretability in a classification task and fusion of metadata and medical imaging data 22 introduced the “gradient-weighted class activation mapping” method, which generates visual explanations by highlighting the regions of the input medical images that are important for the model’s prediction.
For the fusion of multi-omic data and clinical data within the context of prognosis prediction in survival analysis, Jiang et al. 96 and Hao et al. 68 addressed the challenge of the black-box nature of deep neural networks in two different ways. Jiang et al. 96 incorporated Deep SHapley Additive exPlanations (DeepSHAP), a technique that combines deep learning with Shapley values, to explain individual predictions made by the model. The AUTOSurv model, which predicts cancer prognosis, uses DeepSHAP to analyze the SHAP values and quantify how each input feature, such as gene expression or miRNA data, contributes to the final output. This approach enhances the model’s interpretability by linking key pathways with relevant genes, which not only improves the model’s transparency but also highlights potential biomarkers for cancer prognosis.
Similarly, Hao et al. 68 developed CoxPSANet, a deep learning model for cancer survival prediction, which integrates gene expression data at the individual gene level and biological pathways. By establishing sparse connections between the gene and pathway layers, CoxPSANet enhances interpretability by focusing on key biological components and pathways that are relevant to cancer prognosis. For the fusion of imaging and non-imaging data, Xin et al. 60 proposed the “Interpretable Deep Multimodal Fusion” framework, which utilizes Canonical Correlation Analysis to interpret the complex, non-linear associations between modalities. They introduced a cross-modal association (CA) score to assess the impact of each feature, further enhancing the interpretability of the multimodal fusion process.
2.1.4 Imbalance in dimensionality
The term “imbalance in dimensionality” in the context of multimodal data fusion refers to situations in which the different modalities have significantly different numbers of dimensions or features. An imaging modality, for example, could include millions of pixels, but a genomic modality may have only a few hundred or thousand characteristics.
In data fusion, the dimensional imbalance can be problematic since it might lead to an over-reliance on one modality at the expense of others. In some situations, the high-dimensional modality may dominate the fusion process, resulting in poor low-dimensional modality performance. This can lead to information loss and reduced accuracy in final predictions. Several data fusion strategies have been developed to address the challenge of dimensional imbalance. One approach is to use intermediate fusion strategies, in which the modalities are fused gradually rather than completely at once. This can help to reduce the dimensional imbalance by forcing the marginal representations to be of similar size. Another option is to lower the dimensionality of the high-dimensional modality using feature selection or dimensionality reduction algorithms. This can help to prevent the high-dimensional modality from dominating the fusion process and increasing the low-dimensional modality’s performance.
In our review, 7 studies addressed this challenge, using various techniques. For example, Karaman et al. 39 and Partin et al. 25 employed weighted loss functions to assign greater importance to underrepresented modalities or classes, counteracting the overemphasis on high-dimensional data. Similarly, Mao et al. 58 and Li et al. 22 utilized focal loss functions to concentrate learning on harder-to-classify samples, which helped reduce bias toward dominant modalities and enhanced model robustness. Other approaches included non-linear dimensionality reduction, as demonstrated by Madabhushi et al. 80, which aimed to preserve the intrinsic structure of the data while minimizing dimensionality-related imbalances. In a different vein, Xu et al. 49 leveraged recurrent neural networks to model temporal dependencies, helping to better align and integrate multimodal inputs over time.
Although dimensionality imbalance focuses on feature-level disparities, related issues such as class imbalance, particularly common in high-dimensional spaces, can further complicate multimodal classification tasks. Tran and Liatsis 75 addressed this by introducing the RABOC algorithm, a hybrid method combining one-class classification and Real AdaBoost. One-class classification is inherently suited to extreme class imbalance, as it defines decision boundaries using only the target class, while Real AdaBoost boosts performance by focusing on misclassified samples through weighted re-training. Together, these methods effectively handle severe class imbalance in high-dimensional multimodal biometric authentication tasks without overfitting. Similarly, Li et al. 22 tackled class imbalance in multimodal classification through the Deep Multimodal Generative Adversarial Network (DMGAN). This approach generates synthetic features for minority classes via adversarial training, thereby rebalancing the dataset. DMGAN outperforms traditional oversampling techniques like SMOTE (Synthetic Minority Over-sampling Technique) 97, XGBoost (eXtreme Gradient Boosting) 98, Unimodal GAN frameworks 99, and Random Forest 100, demonstrating the effectiveness of generative approaches in dealing with imbalanced multimodal data.
2.1.5 Optimal fusion technique
To make the most out of multimodal data, an optimal fusion strategy is required. The optimal fusion strategy in modeling multimodal data depends on various factors such as the specific task requirements, the characteristics of the modalities involved (e.g., the number of modalities, their relationships, dimensionality, and sampling rates), the available computational resources, trade-offs between model complexity, interpretability, performance, and the ease of implementation 101. Additionally, the selection of an optimal fusion strategy should also consider the quality and reliability of each modality’s data, as well as the level of correlation between the modalities 30. By considering all these factors, the most effective fusion strategy can be determined for specific multimodal data modeling tasks. Overall, an optimal fusion strategy in modeling multimodal data involves integrating and combining multiple modalities in a way that maximizes the benefits of each modality while minimizing the biases and limitations inherent in individual modalities 5.
In addressing the challenge of finding the optimal fusion strategy, many studies in the literature explore different approaches. However, the choice can vary greatly depending on the specific problem. Ramachandram and Taylor 101 reviewed various options for learning an optimal architecture in the context of deep multimodal learning. They highlighted that determining the best fusion strategy depends on two architectural design choices. The first choice relates to the stage at which different modalities should be combined, which can be guided by considering the challenges and advantages of early, late, and intermediate fusion approaches. The second architectural design choice relates to which modalities to combine. They discussed the idea that including all available modalities may not always improve machine learning algorithm performance, suggesting the need for feature selection or optimization of fusion depth. They explored techniques for automatically learning the best ordering and depth of fusion during training to identify the most effective modalities to combine. These techniques include stochastic regularization, treating architecture optimization as a hyperparameter optimization problem using methods like Bayesian optimization, and employing incremental online reinforcement learning. A total of 13 studies in our review addressed this challenge. For example, Xu et al. 102 tackled this challenge by extending the neural architecture search 103, 104, which was originally designed for unimodal data, to incorporate the joint optimization of fusion strategies and modality-specific architectures for multimodal data. This extension allows for searching for unique architectures for distinct modalities and identifies the best strategies to fuse these architectures at the right representation level, and then they utilize state-of-the-art neural architecture search methods to explore and optimize fusion strategies for multimodal data. By jointly optimizing fusion strategies and modality-specific architectures, their approach aims to enhance prediction performance on EHR tasks while controlling for computation costs.
In the fusion of imaging and clinical data, Wang et al. 30 pointed out a common heuristic approach to data fusion in many studies, where data from multiple sources are simply concatenated without proper preprocessing or assessment of fusion quality. This failure to interpret the fused feature space can result in suboptimal model performance, particularly in medical imaging problems where the characteristics of different data sources can vary significantly. Therefore, they proposed modeling the fully connected layer of a deep neural network as a potential function following the classical Gibbs measure. By considering the features of the fully connected layer as random variables governed by state functions representing different data sources, the method quantifies source bias and aims to optimize data fusion quality. The use of a vector-growing encoding scheme, such as positional encoding, helps the conversion of low-dimensional clinical features into a rich feature space to complement high-dimensional imaging features, ultimately enhancing the fusion process in multi-source deep learning problems.
Several studies have highlighted that different modalities often have unequal importance and contributions, necessitating innovative methods to adaptively weight modalities during the fusion process. For instance, Dai et al. 19 employs dual encoders and joint attention mechanisms to assign greater importance to features that are most relevant for Alzheimer’s disease classification. Similarly, Hatami et al. 38 introduces a weighted ensemble method for predicting stroke outcomes, where weights are assigned to imaging module predictions based on their alignment with clinical metadata, thereby improving prediction accuracy. In another approach, Rahaman et al. 26 uses attention mechanisms to prioritize the most informative modalities, mitigating performance degradation that arises from forcing equal contributions. Lastly, Xia et al. 31 incorporates an adaptive fusion layer to dynamically assign modality-specific weights, optimizing their contributions in clustering tasks. Collectively, these studies underscore the critical role of adaptive weighting techniques in enhancing the effectiveness of multimodal fusion by accounting for the varying importance of different modalities. In the classification task, there are various approaches to determining the optimal fusion technique. For example, Pawłowski et al. 5 emphasized the significance of selecting the most appropriate fusion technique. Specifically in classification tasks, they outlined three criteria for making design decisions when selecting the most suitable data fusion technique. These criteria involve evaluating the influence of each modality on the machine learning problem, considering the task type, and assessing memory usage during both training and prediction phases.
There are also two similar works by Khan et al. 16 and Jabeen et al. 15 which determine the optimal fusion technique in two steps. Firstly, they select the best features using the differential evolution selected features. Khan et al. 16 introduced a probability-based fusion method that involves computing probabilities for selected features and then choosing one feature based on the highest probability value. This selected feature is used as a reference for comparison, and features from different vectors are merged into a single matrix. This comparison step helps address the issue of redundant features in the vectors. On the other hand, Jabeen et al. 15 proposed the maximum correlation-based fusion approach, aiming to combine optimal feature vectors by selecting features with strong correlations, discarding those with weak correlations, and merging them into a single vector. By prioritizing features with high correlation coefficients, this fusion method enhances the representational power of the fused vector. This approach ensures that the combined features effectively capture complementary information, leading to improved classification accuracy.
2.2 Other challenges
Beyond the main challenges detailed in this manuscript, there is a range of unique additional issues that do not align with predefined categories yet are still critical for understanding and improving multimodal data integration. In LABEL:Tab:challenges, the specific challenges for “Other Challenges” are explicitly stated alongside their corresponding solutions, as they vary across different studies. One major challenge is extracting high-order interactions within and among data modalities, which is crucial for identifying complex, non-linear relationships that may exist between features. These interactions are particularly relevant in clinical contexts where modalities such as genomics and imaging provide complementary insights. Similarly, capturing the intricate relationships and temporal dynamics across modalities, while accounting for irregularities in patient visit timings, is a significant hurdle. Patient data often exhibits asynchronous patterns, with varying time gaps between visits, making it essential to develop models capable of handling such inconsistencies. High dimensionality adds another layer of complexity, as many modalities, such as imaging and genomics, have a vast number of features. This can lead to intensive computational requirements and the risk of overfitting, necessitating robust dimensionality reduction and feature selection techniques. Capturing intermodality interactions further complicates modeling, as it requires mechanisms that can integrate and prioritize information from heterogeneous sources without losing critical cross-modal correlations. Lastly, heterogeneity in data, whether in terms of formats, scales, or quality, presents a persistent challenge, particularly in ensuring that diverse modalities are integrated cohesively. These challenges highlight the need for innovative approaches, such as advanced attention mechanisms, hybrid fusion strategies, and efficient computational frameworks, to enable robust and interpretable multimodal data integration in clinical applications.
3 Discussion
In this systematic review of multimodal data fusion challenges and solutions in the clinical domains, we have identified and examined various obstacles faced when integrating data from diverse sources such as electronic health records, imaging, wearable devices, and more. Major obstacles were missing modality, limited datasets, interpretability concerns, imbalanced dimensionality, and selecting suitable fusion algorithms. The issue of small data continues to be a critical barrier for effective multimodal fusion, especially in medical settings where large datasets are often unavailable.
One notable aspect of this review is the variety of innovative approaches and methodologies proposed by researchers to address these challenges. For instance, methods such as multitasking learning, matrix completion, variational autoencoders, generative adversarial networks, hybrid graph convolutional networks, and knowledge distillation have been introduced to tackle missing of an entire modality and improve imputation accuracy. Additionally, techniques like data augmentation, transfer learning, and knowledge distillation have been effective in mitigating the impact of small datasets on model performance. However, their application in medical settings still remains constrained. Such methods often rely on access to high-quality pre-trained models or large, diverse datasets, which may not be readily available in the medical context. Furthermore, techniques like data augmentation and feature selection may struggle to fully capture the complex relationships between modalities when data is limited or heterogeneous. Interpretability, a critical aspect in medical applications, has been addressed through attention mechanisms, gradient-based attribution algorithms, and the incorporation of deep learning models with explanatory techniques like DeepSHAP. Moreover, strategies such as feature selection and dimensionality reduction have been proposed to combat the challenge of imbalance in dimensionality, ensuring fair representation across modalities. Determining the optimal fusion technique remains a key consideration, with early, intermediate, and late fusion strategies each offering advantages and drawbacks. Recent advancements in neural architecture search, weighted ensemble methods, and using attention mechanisms hold promise for optimizing fusion strategies and improving model performance.
This review was motivated by the potential of multimodal data modeling for chronic diseases such as COPD and asthma. However, we found a surprising lack of studies specifically addressing the challenges in integrating multimodal data for these conditions. Despite the increasing interest in leveraging diverse data sources, including omics, wearable devices, environmental data, and imaging, to improve the identification of endo- and phenotypes as well as disease progression classification, our systematic search yielded very few relevant studies. The few publications that did mention these diseases either did not focus on modeling challenges or met our exclusion criteria due to limited methodological insights. This gap highlights the need for future research to explore how multimodal integration can enhance disease management for COPD and asthma.
It is important to acknowledge certain limitations in this study. Firstly, the scope of the review may not encompass all existing methodologies and approaches in the field, and there may be additional strategies that were not included in our analysis. Moreover, the effectiveness of the solutions discussed herein may vary depending on the specific characteristics of the dataset and the complexity of the medical problem being addressed. Furthermore, while efforts have been made to select and present the most relevant research findings, there may be biases inherent in the selection process. Additionally, since the systematic search was conducted in October 2023, the rapidly evolving nature of the field means that new advancements may have emerged since the completion of this study.
Future research should focus on specifically developing more tailored methods for handling small datasets in clinical contexts. This could involve incorporating domain-specific knowledge, improving data simulation techniques, or creating hybrid models that integrate both data-driven approaches and clinical insights. By overcoming the small data challenge, we can unlock the full potential of multimodal fusion models, enabling better decision-making and improving healthcare delivery across a variety of clinical applications.
4 Methods
4.1 Search strategy
To systematically identify relevant articles, a structured search strategy was employed (see Table 4). The search was conducted using two major databases, PubMed and manual search from Google Scholar. The time frame for the search was limited until October 2023 for PubMed (search date: 2023/10/09) and until end of 2023 for Google Scholar (search date: 2024/01/02). The strategy was designed to identify studies addressing challenges and methodologies in modeling multimodal data, with a focus on clinical applications. First, studies were required to explicitly mention “model fusion”, “data fusion,” or “multimodal” in their title, ensuring relevance to the topic of multimodal data modeling. To ensure comprehensive coverage, the Boolean operator OR was used within each keyword group to combine related terms. Next, the search included terms related to key research purposes, such as “disease prediction”, “diagnosis”, “clinical outcome prediction”, “modeling”, or “exacerbation”, to capture studies with clinical applications. Additionally, the search incorporated terms representing diverse data modalities, including “electronic health records”, “imaging data”, “wearable devices”, “longitudinal data”, “temporal data”, or “time-series data”, to ensure coverage of a wide range of multimodal datasets. To further refine the results, the term “challenges” was included to focus on studies discussing methodological or analytical obstacles related to multimodal data integration. Finally, the Boolean operator AND was used to combine these components. The final search query combined these elements to ensure a comprehensive yet focused selection of studies. In addition, a manual search was performed to identify any potentially relevant studies not captured in the database search.
4.2 Inclusion and exclusion criteria
Two authors, MF and MW, screened the titles and abstracts of all identified studies to determine their relevance according to prespecified inclusion and exclusion criteria. We selected studies that specifically addressed challenges in the modeling of multimodal data. Study publications were included if they focused on methodologies, approaches, or frameworks for integrating, fusing, or analyzing data from multiple modalities. Studies were excluded if they focused exclusively on multimodal imaging or utilized only audio and video data, as the review aimed to encompass a broader range of multimodal data types. Review articles were excluded to maintain an emphasis on original research. We also excluded publications addressing clinical management, therapeutic interventions, or non-technical challenges (e.g., treatment modalities, surgical techniques, or clinical guidelines) without discussing modeling challenges in multimodal data fusion. Similarly, studies focusing on a single data modality (e.g., imaging, clinical data, or sensor data) without addressing challenges in combining multiple data sources were excluded. Publications centered on clinical applications, such as diagnosis, treatment protocols, or medical decision-making, were included only if they also explored computational modeling or challenges related to multimodal data integration. Additionally, case studies, clinical trial results, and studies describing specific treatments (e.g., surgical procedures, medical devices, or drug interventions) were excluded unless they explicitly addressed how multimodal data were modeled or fused to improve decision-making, diagnosis, or treatment. Finally, although “multimodal” was used as a keyword in the search strategy and required to be present in the title, some publications were excluded because their focus was unrelated to data fusion or computational modeling (e.g., multimodal therapy or multimodal plasma), thus falling outside the scope of this review.
4.3 Data extraction
For the data extraction process, we categorized the included publications according to the specific challenges they addressed. We identified five main technical challenges in modeling multimodal data: (i) missing entire modalities and incomplete datasets, (ii) small data (data sparsity), (iii) interpretability, (iv) imbalance in dimensionality, and (v) identifying the optimal fusion technique. We extracted information on the specific challenges each study tackled within the broader scope of the five identified challenges. We also captured the proposed solutions for overcoming these challenges. We further specifically collected information about the fusion strategies employed in each study. The choice of fusion technique directly relates to the challenge of finding the optimal fusion method (challenge v). Some strategies may also help mitigate other challenges, such as handling incomplete datasets (challenge i) or balancing dimensionality (challenge iv) by integrating information from multiple sources effectively. Finally, because different medical settings encounter distinct modeling challenges, we recorded the clinical domain of each study to link it with the challenges addressed. For example, missing modalities are particularly frequent in fields like neurology and oncology, reflecting the difficulty of integrating heterogeneous sources (such as imaging, genomics, and clinical notes) in these complex conditions Small datasets also occur more often in these fields because rare conditions or challenging data collection keep sample sizes low.
Five co-authors conducted the screening process. Selected full-text studies were distributed among MF, MW, MB, ChG, and JK for the eligibility check and extracting relevant data according to predefined criteria. After the independent evaluations, the results were cross-checked, and a final decision was made on which studies to include in the systematic review.
Data availability
All data supporting the findings of this study are available within the paper and its Supplementary Information files.
Acknowledgements
The work of MF, SW, NH, HB, and NB has been funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 499552394 – SFB 1597. The work of MW, ChG, MB, JK, and HB was partially funded by the Federal Ministry of Education and Research (BMBF) – CALM-QE project, under grant numbers 01ZZ2318B, 01ZZ2318L, and 01ZZ2318I.
Authors’ contributions
MF, MW, HB, and NB contributed to the study’s conception and design. MF, MW, MB, JK, and ChG were involved in the systematic search process. NH extracted the medical domain and disease type of the articles. MF extracted the data, while MW reviewed the extraction. MF wrote the initial draft of the manuscript, with NB providing supervision throughout the process. All authors provided critical feedback and approved the final version.
Competing interests
The authors declare that they have no competing interests.
Appendix A .
Table 4: PubMed Search Query 2023/10/09
Number Query Results
5 #1 AND #2 AND #3 AND #4 801
4 challenges 1,193,213
3 Electronic Health records OR Images OR Imaging data OR Wearables OR longitudinal OR temporal OR time series 5,304,766
2 disease prediction OR diagnosis OR clinical outcome prediction OR modeling OR exacerbation 15,297,673
1 model fusion[Title] OR data fusion[Title] OR multimodal[Title] 18,205